OpenStudy (anonymous):

find the value. logarithm of 125/8 to the base 2.5

OpenStudy (anonymous):

Well you have: \[\log_{5/2}(125/8)=\log_{5/2}((\frac{5}{2})^3)=3\].

OpenStudy (anonymous):

thankx

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

\[(\log_{10} 125/8)/(\log_{10}2.5) \] now just do this on your calculator

OpenStudy (anonymous):

once you are satisfied with an answer you can select 'good answer'